OpenStudy (anonymous):

f(x)=x-4 and g(x)=x^2. Find (gxf) (-1)

OpenStudy (anonymous):

Please help me!

OpenStudy (cruffo):

Can you please clarify: what does fxg mean? Do you mean f times g?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

f times g

OpenStudy (cruffo):

There are actually two ways to do this problem.

OpenStudy (anonymous):

May you please show me both.

OpenStudy (cruffo):

Since you are given a number for x, that is x = -1 in this problem, you can find the individual function values and multiply them. Find f(-1): Find g(-1):

OpenStudy (anonymous):

oh! Thank you! After finding both, using x, you multiply them together?

OpenStudy (cruffo):

yep. What did you get?

OpenStudy (anonymous):

Yes! Thank-You so much!!!!

OpenStudy (cruffo):

cool. did you want to see the other way - it involves a bit more algebra...

OpenStudy (anonymous):

yes please.

OpenStudy (cruffo):

The other way is to multiply the two functions, simplify, then plug in -1 for x: First: \[(fg)(x) = (x - 4) * x^2 = x^3 - 4x^2 \] then \[ (fg)(-1) = (-1)^3 - 4(-1)^2\] Should work out to be the same answer.

OpenStudy (anonymous):

Okay, I understand both ways! Thank-you!